# VidCtx: Context-aware Video Question Answering with Image Models

### v1

@@ Abstract
To address computational and memory limitations of Large Multimodal Models in the Video Question-Answering task, several recent methods extract textual representations per frame (e.g., by captioning) and feed them to a Large Language Model (LLM) that processes them to produce the final response. However, in this way, the LLM does not have access to visual information and often has to process repetitive textual descriptions of nearby frames. To address those shortcomings, in this paper, we introduce VidCtx, a novel training-free VideoQA framework which integrates both modalities, i.e. both visual information from input frames and textual descriptions of others frames that give the appropriate context. More specifically, in the proposed framework a pre-trained Large Multimodal Model (LMM) is prompted to extract at regular intervals, question-aware textual descriptions (captions) of video frames. Those will be used as context when the same LMM will be prompted to answer the question at hand given as input a) a certain frame, b) the question and c) the context/caption of an appropriate frame. To avoid redundant information, we chose as context the descriptions of distant frames. Finally, a simple yet effective max pooling mechanism is used to aggregate the frame-level decisions. This methodology enables the model to focus on the relevant segments of the video and scale to a high number of frames. Experiments show that VidCtx achieves competitive performance among approaches that rely on open models on three public Video QA benchmarks, NExT-QA, IntentQA and STAR.
@@ I Introduction
Video-Language modeling is an inherently resource-intensive task due to the necessity of processing a large number of frames, in order to generate accurate representations of the contents of a video. Recently, a considerable number of Video Large Language Models have emerged [1] [2] [3] [4] [5], which combine a visual encoder with a Large Language Model (LLM). These models have achieved state-of-the-art performance on a wide variety of video understanding tasks. However, they are typically limited by the maximum number of input tokens, and therefore frames, that they can process in one forward pass.
In order to tackle these problems, prior works have proposed sophisticated frame selection policies [6] [7] [8], with the goal of selecting the most salient frames of the input video. In a different direction, several works [9] [10] [11] have proposed extracting textual information in the form of captions from the input video and using an LLM to process the extracted captions. These approaches rely on the ability of LLMs to reason over long input prompts and produce accurate results. However, this has been proven to be a challenging task [12] [13]. Furthermore, by converting the input video to a textual representation the model no longer has access to the rich visual information contained in the video; this reduces the final performance of question answering.
@@
VidCtx [^1] Code will be publicly released upon acceptance. is not tailored to a specific pre-trained LMM, and requires no additional training or fine-tuning. Furthermore, it can process any number of frames since it is not constrained by the maximum context length of the language model. Our experimental results on three publicly available Video Question Answering benchmarks show that our VidCtx framework achieves competitive performance among approaches that rely on comparable open models.
@@ II-A Large Language Models for Video
The combination of a Large Language Model with a visual encoder or adapter module has resulted in a wide variety of models [1] [2] [3] [4] [5] [19] that have achieved state-of-the-art performance on several video understanding tasks. These models primarily benefit from large-scale pre-training and instruction tuning, therefore, their performance is strongly dependent upon the quantity and quality of the available training data.
@@ III-B Question-aware Caption Extraction
* VFC [27] pre-train their model on 0.5M video-caption pairs. SeViLA [6] trains a localizer component on 10K videos. InternVideo [28] and VideoChat2 [1] utilize large-scale pre-training on video datasets with over 10M video-caption pairs.
@@ III-C Context-aware Video Question Answering
where $d_{i}$ is the $i$-th decision, $r(i)$ is the index of the distant frame and $I_{vqa}$ is the concatenation of the instruction prompt and the given question $Q$. We choose to share information between frames that are exactly half the duration of the video apart, in order to ensure the greatest diversity between the contextual captions and the considered. We also insert the relevant temporal specifier (“earlier” or “later”), in order to inform the language model about the sequence of events in the video.
@@ IV-B Datasets and Metrics
    IntentQA [25] is a video QA dataset, derived from NExT-QA, that focuses on intent reasoning. We perform our experiments on the test set that contains 2.1K question-answer pairs.
@@
    STAR [26] is a video QA dataset that contains 60K situated questions, divided into 4 categories (Interaction, Prediction, Sequence, Feasibility). Similarly to the relevant literature, we perform our experiments on the validation set, which contains 7K questions.
@@ V Conclusion
Acknowledgment
This work was supported by the EU Horizon Europe programme under grant agreement 101070109 TransMIXR.

### v2

@@ Abstract
To address computational and memory limitations of Large Multimodal Models in the Video Question-Answering task, several recent methods extract textual representations per frame (e.g., by captioning) and feed them to a Large Language Model (LLM) that processes them to produce the final response. However, in this way, the LLM does not have access to visual information and often has to process repetitive textual descriptions of nearby frames. To address those shortcomings, in this paper, we introduce VidCtx, a novel training-free VideoQA framework which integrates both modalities, i.e. both visual information from input frames and textual descriptions of others frames that give the appropriate context. More specifically, in the proposed framework a pre-trained Large Multimodal Model (LMM) is prompted to extract at regular intervals, question-aware textual descriptions (captions) of video frames. Those will be used as context when the same LMM will be prompted to answer the question at hand given as input a) a certain frame, b) the question and c) the context/caption of an appropriate frame. To avoid redundant information, we chose as context the descriptions of distant frames. Finally, a simple yet effective max pooling mechanism is used to aggregate the frame-level decisions. This methodology enables the model to focus on the relevant segments of the video and scale to a high number of frames. Experiments show that VidCtx achieves competitive performance among approaches that rely on open models on three public Video QA benchmarks, NExT-QA, IntentQA and STAR. Our code is available at https://github.com/IDT-ITI/VidCtx.
@@ I Introduction
Video-Language modeling is an inherently resource-intensive task due to the necessity of processing a large number of frames, in order to generate accurate representations of the contents of a video. Recently, a considerable number of Video Large Language Models have emerged [1] [2] [3] [4] [5] [6] [7], which combine a visual encoder with a Large Language Model (LLM). These models have achieved state-of-the-art performance on a wide variety of video understanding tasks. They are pre-trained on large video corpora, thus they are typically not considered zero-shot approaches. Furthermore, they are generally limited by the maximum number of input tokens, and therefore frames, that they can process in one pass.
In order to tackle this problem, prior works have proposed sophisticated frame selection policies [8] [9] [10], with the goal of selecting the most salient frames of the input video. In a different direction, several works [11] [12] [13] have proposed extracting textual information in the form of captions from the input video and using an LLM to process the extracted captions. These approaches rely on the ability of LLMs to reason over long input prompts and produce accurate results. However, this has been proven to be a challenging task [14] [15]. Furthermore, by converting the input video to a textual representation the model no longer has access to the rich visual information contained in the video; this reduces the final performance of question answering.
@@
VidCtx is not tailored to a specific pre-trained LMM, and requires no additional training or fine-tuning. Furthermore, it can process any number of frames since it is not constrained by the maximum context length of the language model. Our experimental results on three publicly available Video Question Answering benchmarks show that our VidCtx framework achieves competitive performance among approaches that rely on comparable open models.
@@ II-A Large Language Models for Video
The combination of a Large Language Model with a visual encoder or adapter module has resulted in a wide variety of models [1] [2] [3] [4] [5] [6] [7] that have achieved state-of-the-art performance on several video understanding tasks. These models primarily benefit from large-scale pre-training and instruction tuning, including training on video corpora, therefore their performance is strongly dependent upon the quantity and quality of the available training data.
@@ III-B Question-aware Caption Extraction
* VFC [28] is pre-trained on 0.5M video-caption pairs. SeViLA [8] contains a localizer component trained on 10K videos. InternVideo [29] and VideoChat2 [5] utilize large-scale pre-training on video datasets with over 10M video-caption pairs.
@@ III-C Context-aware Video Question Answering
where $d_{i}$ is the $i$-th decision, $r(i)$ is the index of the distant frame and $I_{vqa}$ is the concatenation of the instruction prompt and the given question $Q$. We choose to share information between frames that are exactly half the duration of the video apart, in order to ensure the greatest diversity between the contextual captions and the considered frame. We also insert the relevant temporal specifier (“earlier” or “later”), in order to inform the language model about the sequence of events in the video.
@@ IV-B Datasets and Metrics
    IntentQA [26] is a video QA dataset, derived from NExT-QA, that focuses on intent reasoning, with an average video duration of 44 seconds. We perform our experiments on the test set that contains 2.1K question-answer pairs.
@@
    STAR [27] is a video QA dataset that contains 60K situated questions, divided into 4 categories (Interaction, Prediction, Sequence, Feasibility), with an average video duration of 12 seconds. Similarly to the relevant literature, we perform our experiments on the validation set, which contains 7K questions.